Question Number 3: A new drug is proposed to lower total cholesterol and a study is designed to evaluate the efficacy of the drug in lowering cholesterol. Fifteen patients agree to participate in the study and each is asked to take the new drug for 6 weeks. However, before starting the treatment, each patient's total cholesterol level is measured. The initial measurement is a pre-treatment or baseline value. After taking the drug for 6 weeks, each patient's total cholesterol level is measured again and the data are shown below. Subject Identification Number   Baseline   6 Weeks 1   215   205 2   190   156 3   230   190 4   220   180 5   214   201 6   240   227 7   210   197 8   193   173 9   210   204 10   230   217 11   180   142 12   260   262 13   210   207 14   190   184 15   200   193 Is there statistical evidence of a reduction in mean total cholesterol in patients after using the new medication for 6 weeks?
Answer #1

First of all note that the data has been collected on the same set of people before and after treatment with new drug. As the data has been collected using repeated samples, so the data is paired and thus we would use a paired sample t-test.

As we want to test if the drug has reduced the mean cholestrol in patients, so we would want the mean of differences for population to be positive such that the data before treatment are higher than post treatment.

Define 11 and 112 as the population mean cholestrol level before and after administration of drug into the body respectively. And let D be the difference -

HD 11 H2

So, the hypothesis would be lower tailed written as -

Null Hypothesis - H0: \mu_{D} = 0

Alternate Hypothesis - H1: \mu_{D} \neq 0

Now, create a separate column for difference between the values of 'Baseline' and '6 Weeks After' as shown -

Identification Number Baseline (X1i) 6 Weeks after (X2i) Difference (di)
1 215 205 10
2 190 156 34
3 230 190 40
4 220 180 40
5 214 201 13
6 240 227 13
7 210 197 13
8 193 173 20
9 210 204 6
10 230 217 13
11 180 142 38
12 260 262 -2
13 210 207 3
14 190 184 6
15 200 193 7

Now, you can either use excel to directly get the test result or you can do it manually.

1) Using Excel

Use excel's 'Data Analysis' add-on under the 'Data' tab to access the 't-Test: Paired Two Sample for Means' as shown -

File Home Insert Draw Page Layout Formulas Data Review View Help Power Pivot Share From Text/CSV Queries & Connections I UN R

Then in the next prompt box, enter the variable ranges as the 'Baseline' and '6 Weeks Later' data. As the hypothesized mean difference is '0', so put the same in prompt. Also make sure to mark the 'Labels' option as we have label in our first row. Your input should look similar to this -

J1 X E S t-Test: Paired Two Sample for Means ? X 1 2 ܢܙ Input Variable 1 Range: OK 3 $G$1:$G$16 1 4 Variable 2 Range: Cancel

Note that you can set level of significance as per your needs. By default its chosen as 0.05.

This should give you following result -

t-Test: Paired Two Sample for Means
Baseline 6 Weeks after
Mean 212.8 195.8666667
Variance 450.8857143 824.2666667
Observations 15 15
Pearson Correlation 0.881282389
Hypothesized Mean Difference 0
df 14
t Stat 4.630004243
P(T<=t) one-tail 0.000194797
t Critical one-tail 1.761310136
P(T<=t) two-tail 0.000389595
t Critical two-tail 2.144786688

As we had a one-tailed test, so note that the p-value for a one tailed test = 0.000194797 \approx 0.0002

As the p-value is less than the significance level, so we reject the null hypothesis and conclude that there is enough evidence to support the claim that the cholestrol level has decreased 6 weeks after administration of drug in the body.

__________________________________________________

Solving it manually -

Calculate the sample mean differences and standard deviation of differences using formula -

\overline{X_{D}} = \frac{\sum d_{i}}{n}

s_{D} = \sqrt{\frac{\sum (d_{i} - \overline{X_{D}})}{n-1}}

You can calculate the values as -

Identification Number Baseline (X1i) 6 Weeks after (X2i) Difference (di) d_{i} - \overline{X_{D}} (d_{i} - \overline{X_{D}})^2
1 215 205 10 -6.93333 48.07111
2 190 156 34 17.06667 291.2711
3 230 190 40 23.06667 532.0711
4 220 180 40 23.06667 532.0711
5 214 201 13 -3.93333 15.47111
6 240 227 13 -3.93333 15.47111
7 210 197 13 -3.93333 15.47111
8 193 173 20 3.066667 9.404444
9 210 204 6 -10.9333 119.5378
10 230 217 13 -3.93333 15.47111
11 180 142 38 21.06667 443.8044
12 260 262 -2 -18.9333 358.4711
13 210 207 3 -13.9333 194.1378
14 190 184 6 -10.9333 119.5378
15 200 193 7 -9.93333 98.67111
Sum 254 2808.933

\overline{X_{D}} = \frac{\sum d_{i}}{n} = \frac{254}{15} = 16.9333

s_{D} = \sqrt{\frac{2808.933}{14}} = 14.164668

Then the test statistic is -

t = \frac{\overline{X_{D}} - \mu_{D}}{\frac{s_{D}}{\sqrt{n}}}

So, we get -  

t = \frac{16.9333 - 0}{\frac{14.16468}{\sqrt{15}}} = 4.630

The degree of freedom of test = n-1 = 15 - 1 = 14

Significance level = \alpha = 0.05

So, critical value of test statistic = t0.05,14 = 1.761

As calculated value of test statistic = 4.63 > critical value of test statistic = 1.761, so we reject the null hypothesis and conclude that there is enough evidence in the data to support the claim that the mean cholestrol of population decreases after 6 weeks of administration of drug into body.

____________________________________________

Please ask if you have any doubt(s) in comment section.

____________________________________________________________________